OpenStudy (anonymous):

distance between points (2, 8) and (7, 5)

OpenStudy (adamaero):

use the distance formula

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

i forgot

OpenStudy (anonymous):

@adamaero

OpenStudy (adamaero):

(I'm tagged automatically because I replied) http://www.purplemath.com/modules/distform.htm

OpenStudy (anonymous):

i got 4

OpenStudy (anonymous):

5.8?

OpenStudy (anonymous):

hurry

OpenStudy (anonymous):

Yes, the distance formula is \(\color{blue}{d=\sqrt{(x_2 - x_1)^2+(y_2-y_1)^2}}\) x2 = 7 x1 = 2 y2 = 5 y1 = 8 PLug them in \(\color{blue}{d=\sqrt{(7 - 2)^2+(5-8)^2}}\) Now

OpenStudy (anonymous):

What is 7 - 2 and 5 - 8?

OpenStudy (anonymous):

is the answer 5.8

OpenStudy (anonymous):

25+9 34 34 squared 5.8

OpenStudy (anonymous):

its not multiple choice

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

THe answer is rounded to 5.8

OpenStudy (anonymous):

Hold on a sec

OpenStudy (anonymous):

Ok yeah

OpenStudy (anonymous):

5.83

OpenStudy (anonymous):

7.21 units

OpenStudy (anonymous):

google told me